write an equation of the line that passes through the given point and has the given slope.
(2,3),m=1/2

Answer by checkley71(8403)

Y=mX+b WHERE m=SLOPE & b=Y INTERCEPT.
3=1/2*2+b
3=1+b
b=3-1
b=2 THUS THE LINE EQUATION IS:
Y=X/2+2 ANSWER.